Operating Systems and Systems Integration

Setting up the latest Mozilla, with the Java plugin, Flash 6 and Acrobat Plugin

Contents

1 Background  1
1.1 The Automounter . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  1
2 Procedure  1
2.1 Installing Mozilla . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  1
2.2 Installing the Java SDK . . . . . . . . . . . . . . . . . . . . . . . . .  2
2.3 Adding Macromedia Flash 6 . . . . . . . . . . . . . . . . . . . . . . .  3
2.4 Installing the Acrobat reader 5.0.6 . . . . . . . . . . . . . . . . . . . .  3
2.5 Testing Mozilla . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  5
3 Selecting useful options in Mozilla  5
3.1 Tabbed Browsing . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  5
3.2 Some other Miscellaneous Settings . . . . . . . . . . . . . . . . . . . .  5
3.3 Additional features Mozilla has that IE Lacks . . . . . . . . . . . . .  6

1 Background

The current version of Mozilla (1.2.1, on 12 December 2002) is a useful improvement over the one that comes with Red Hat 8.0. Here I give a quick guide on how to install it, set up the Acrobat Reader, the jre plugin, and install the Macromedia Flash 6 plugin.

I will also explain some of the good features of Mozilla.

1.1 The Automounter

There is a program called the automounter which automatically mounts local and network file systems. The setup is via our ldap directory service, and it includes the ftp site on our server ictlab, which is accessible on your computer when you access the directory

/home/nfs. So for example, the redhat-8.0 directory is at /home/nfs/redhat-8.0, and
Red Hat updates are available at /home/nfs/redhat-8.0/updates.

2 Procedure

2.1 Installing Mozilla

We have the Mozilla software packages on our server ictlab, but you can download them

from http://www.mozilla.org/releases/ if you are at home.

Setting up the latest Mozilla, with the Java plugin, Flash 6 and Acrobat Plugin Operating Systems and Systems Integration

2

1. Change to the mozilla directory:

$ cd /home/nfs/mozilla

2. and install mozilla:

$ sudo rpm -Uhv *.rpm

2.2 Installing the Java SDK

Note that here we will install the Java sdk and Java run time environment, to be used by the browser.

1. Change to the directory on ictlab which contains the jdk:

$ cd /home/nfs/jdk

2. and install it:

$ sudo rpm -Uhv j2sdk-1_4_1_01-fcs-linux-i586.rpm

3. Now include the executables on your PATH by adding these lines (shown below) to your /etc/profile login script:

$ xhost +localhost
$ sudo -v
$ sudo emacs /etc/profile &

Before the line that says:

unset pathmunge

add the lines:

JAVA_HOME=$(echo /usr/java/j2*)
pathmunge $JAVA_HOME/bin after
export JAVA_HOME

4. Finally, add the plugin to Mozilla’s plugin directory. We create a symbolic link (see the handout on symbolic links: we cover this topic soon):

$ cd /usr/lib/mozilla-1.2.1/plugins
$ sudo ln -s \
../../../java/j2sdk1.4.1_01/jre/plugin/i386/ns610/libjavaplugin_oji.so .

Note: I just keep pressing the [figure: tab]key; I do not memorise this file name!

Setting up the latest Mozilla, with the Java plugin, Flash 6 and Acrobat Plugin Operating Systems and Systems Integration

3

2.3 Adding Macromedia Flash 6

If you read the release notes for Mozilla 1.2.1, it says that we should really use Macromedia Flash version 6, not the current version 5. You can find a link to the download site for

Flash 6 from the Mozilla release notes, but I reproduce it here: http://www.macromedia.com/software/flashplayer/special/beta/.

To install it from our server:

1. Change to your /tmp directory:

     $ cd /tmp

2. unpack it:

     $ tar xvzf /home/nfs/mozilla/install_flash_player_6_linux.tar.gz

3. Then begin the installation:

     $ cd install_flash_player_6_linux
     $ sudo ./flashplayer-installer

4. Accept defaults, except when it asks you where it should be installed (it offers

     /usr/lib/mozilla), say /usr/lib/mozilla-1.2.1

2.4 Installing the Acrobat reader 5.0.6

Note first that Adobe’s Acrobat reader does not work with Unicode; it would not work for me until I set my locale to US English. If that is okay for you, then proceed. I have mixed feelings about this software; sometimes I like it, othertimes I find it does not work very well. I use it nonetheless. You can download the acrobat reader from http://www.adobe.com, or you can install it from our server, ictlab.

1. Change to your /tmp directory:

     $ cd /tmp

2. untar the tarball:

     $ tar xvzf /home/nfs/adobe/linux-506.tar.gz

3. install it, accepting defaults:

     $ sudo ./INSTALL

4. Create a symbolic link to it onto your PATH:

Setting up the latest Mozilla, with the Java plugin, Flash 6 and Acrobat Plugin Operating Systems and Systems Integration

4

$ sudo ln -s /usr/local/Acrobat5/bin/acroread /usr/local/bin

5. Optionally, you can install the Chinese fonts:

$ cd /tmp
$ tar xvzf /home/nfs/adobe/fonts-5.0.6/chtfont.tar.gz
$ cd CHTKIT
$ sudo ./INSTLANG

Accept defaults and you are done.

6. Setting the language to en US: Edit the file /etc/sysconfig/i18n, and change the line:

LANG=en_US.UTF8

to:

LANG=en_US

7. To apply this change permanently, you may need to reboot, but you can apply it to your present session by typing:

$ export LANG=en_US

and try the acrobat reader on some PDF document you have:

$ acroread workshop-notes.pdf

8. To test the plugin in the browser, you need to start mozilla with LANG set to en US:

$ mozilla

After you reboot, you can start Mozilla by clicking on the globe symbol in the panel at the bottom of your screen, and the acrobat reader should work.

Setting up the latest Mozilla, with the Java plugin, Flash 6 and Acrobat Plugin Operating Systems and Systems Integration

5

2.5 Testing Mozilla

Now you are ready to check that everything is set up right:

1. Start Mozilla by clicking on the globe in your panel (the tool bar at the bottom of your screen)

2. Select Help⟶About Plugins

3. Verify that you can see:

   (a) Java(TM) Plug-in 1.4.1 01-b01
   (b) nppdf.so

(c) Flash version 6

4. Verify that the acrobat reader works on web documents.

3 Selecting useful options in Mozilla

Mozilla has some nice features that you should enable. You can also set your home page:

1. Click on Navigator and set your home page to http://nicku.org/ossi/. You should be aware that finally, full screen mode works in Linux. Press [figure: F11]to toggle full screen mode.

3.1 Tabbed Browsing

Tabbed browsing is a nice feature. Each web page opens in the same browser frame, but in another tab. Try it out:

1. Select Edit⟶Preferences

2. Click on the triangle next to Navigator and click on Tabbed Browsing.

3. Click on the two check boxes at the bottom under “Open tabs instead of windows

for” Now when you click on a link with the middle mouse button (or both buttons if you have only two), the web page will open in a new tab.

3.2 Some other Miscellaneous Settings

1. If you are developing web pages, then it is important to increase the checking for changes in the web document. I have the setting under Advanced⟶Cache set to “Compare the page in the cache to the page on the network” to “Every time I view the page”.

2. Under Advanced⟶Proxies, I strongly suggest using Richard Yim’s proxy configuration script. Select “Automatic proxy configuration file” and enter: http://tyintra.vtc.edu.hk/proxy.proxy.

3. Set your search engine to Google: under Navigator⟶Internet Search, select “Search using: Google”.

Nick Urbanik <nicku(at)nicku.org>  ver. 1.4

Setting up the latest Mozilla, with the Java plugin, Flash 6 and Acrobat Plugin Operating Systems and Systems Integration

6

3.3 Additional features Mozilla has that IE Lacks

John Dvorak from PC Magazine makes an interesting comparison of Mozilla with IE

at this article: http://www.pcmag.com/article2/0,4149,715464,00.asp. The main

features he points out are (I wouldn’t know; I use IE for nothing but for applying Windows updates):

• You can expand the font size for any fonts in Mozilla

• You can block pop-up, pop-behind windows (go to Advanced⟶Scripts & Plugins and deselect “Allow scripts to: Open unrequested windows”, among other choices).

• Mozilla is standards compliant. Microsoft don’t bother much with standards in IE.

• Printing works better in Mozilla: it doesn’t cut off the right hand side of the page when you print.